OpenStudy (anonymous):

f(x)=2/x and g(x)= sqrt x 1) f+g 2) g(f(x)) 3) f(g(x))

OpenStudy (anonymous):

for the first one, put a plus sign between \(f\) and \(g\)

OpenStudy (anonymous):

for the second one, start with \[g(f(x))=g(\frac{2}{x})\] then since \(g(\spadesuit)=\sqrt{\spadesuit}\) you have \[g(\frac{2}{x})=\sqrt{\frac{2}{x}}\]